A petrol station has two pumps. The service time follows the exponential distribution with mean 4 minutes and cars arrive for service in a Poisson process at the rate of 10 cars per hour. Find the probability that a customer has to wait for service. What proportion of time the pump remains idle?  


1
Expert's answer
2022-03-23T15:36:43-0400

"\mu=\frac{60}{4}=15"

"\lambda=10"

"P(Customer has to wait for service)=\frac{1}{15-10}=0.2"

Proportion of time the pump stays idle = "\frac{10}{15}=0.66"